\begin{thebibliography}{}

\bibitem[\protect\citename{Bordes \bgroup et al.\egroup
  }2011]{bordes2011learning}
Antoine Bordes, Jason Weston, Ronan Collobert, and Yoshua Bengio.
\newblock 2011.
\newblock Learning structured embeddings of knowledge bases.
\newblock In {\em Twenty-Fifth AAAI Conference on Artificial Intelligence}.

\bibitem[\protect\citename{Bordes \bgroup et al.\egroup
  }2013]{bordes2013translating}
Antoine Bordes, Nicolas Usunier, Alberto Garcia-Duran, Jason Weston, and Oksana
  Yakhnenko.
\newblock 2013.
\newblock Translating embeddings for modeling multi-relational data.
\newblock In {\em Advances in Neural Information Processing Systems}, pages
  2787--2795.

\bibitem[\protect\citename{Budanitsky and Hirst}2006]{BudanitskyHirst:06}
Alexander Budanitsky and Graeme Hirst.
\newblock 2006.
\newblock {Evaluating WordNet-based measures of Lexical Semantic Relatedness}.
\newblock {\em Computational Linguistics}, 32(1):13--47.

\bibitem[\protect\citename{Camacho-Collados and
  Navigli}2016]{camachocolladosrepeval:2016}
Jos{\'e} Camacho-Collados and Roberto Navigli.
\newblock 2016.
\newblock {Find the word that does not belong: A Framework for an Intrinsic
  Evaluation of Word Vector Representations}.
\newblock In {\em Proceedings of the ACL Workshop on Evaluating Vector Space
  Representations for NLP}, Berlin, Germany.

\bibitem[\protect\citename{Camacho-Collados \bgroup et al.\egroup
  }2015a]{camachocolladosetal:2015}
Jos\'e Camacho-Collados, Mohammad~Taher Pilehvar, and Roberto Navigli.
\newblock 2015a.
\newblock {\GG{1}}{N}{A}{S}{A}{R}{I}: a {N}ovel {A}pproach to a
  {S}emantically-{A}ware {R}epresentation of {I}tems.
\newblock In {\em Proceedings of NAACL}, pages 567--577.

\bibitem[\protect\citename{Camacho-Collados \bgroup et al.\egroup
  }2015b]{camachocolladosetal:2015b}
Jos\'{e} Camacho-Collados, Mohammad~Taher Pilehvar, and Roberto Navigli.
\newblock 2015b.
\newblock {{\GG{2}}A Unified Multilingual Semantic Representation of Concepts}.
\newblock In {\em Proceedings of ACL}, pages 741--751, Beijing, China.

\bibitem[\protect\citename{Camacho-Collados \bgroup et al.\egroup
  }2015c]{camacho2015framework}
Jos{\'e} Camacho-Collados, Mohammad~Taher Pilehvar, and Roberto Navigli.
\newblock 2015c.
\newblock {{\GG{3}}A Framework for the Construction of Monolingual and
  Cross-lingual Word Similarity Datasets}.
\newblock In {\em Proceedings of the 53rd Annual Meeting of the Association for
  Computational Linguistics -- Short Papers}, pages 1--7, Beijing, China.

\bibitem[\protect\citename{Camacho-Collados \bgroup et al.\egroup
  }2016a]{camachocollados:lrec2016}
Jos\'{e} Camacho-Collados, Claudio~Delli Bovi, Alessandro Raganato, and Roberto
  Navigli.
\newblock 2016a.
\newblock A large-scale multilingual disambiguation of glosses.
\newblock In {\em Proceedings of the 10th International Conference on Language
  Resources and Evaluation (LREC 2016)}, pages 1701--1708, Portoroz, Slovenia,
  May. European Language Resources Association (ELRA).

\bibitem[\protect\citename{Camacho-Collados \bgroup et al.\egroup
  }2016b]{camachocollados:2016aij}
Jos\'e Camacho-Collados, Mohammad~Taher Pilehvar, and Roberto Navigli.
\newblock 2016b.
\newblock {{NASARI: Integrating Explicit Knowledge and Corpus Statistics for a
  Multilingual Representation of Concepts and Entities}}.
\newblock {\em Artificial Intelligence}.

\bibitem[\protect\citename{Chen \bgroup et al.\egroup }2014]{chenunified:2014}
Xinxiong Chen, Zhiyuan Liu, and Maosong Sun.
\newblock 2014.
\newblock A unified model for word sense representation and disambiguation.
\newblock In {\em Proceedings of EMNLP}, pages 1025--1035, Doha, Qatar.

\bibitem[\protect\citename{Chen \bgroup et al.\egroup }2015]{chen2015improving}
Tao Chen, Ruifeng Xu, Yulan He, and Xuan Wang.
\newblock 2015.
\newblock Improving distributed representation of word sense via wordnet gloss
  composition and context clustering.
\newblock In {\em Proceedings of the 53rd Annual Meeting of the Association for
  Computational Linguistics and the 7th International Joint Conference on
  Natural Language Processing -- Short Papers}, pages 15--20, Beijing, China.

\bibitem[\protect\citename{Crouch}1988]{Crouch:1988}
C.~J. Crouch.
\newblock 1988.
\newblock A cluster-based approach to thesaurus construction.
\newblock In {\em Proceedings of the 11th Annual International ACM SIGIR
  Conference on Research and Development in Information Retrieval}, SIGIR '88,
  pages 309--320.

\bibitem[\protect\citename{Curran and Moens}2002]{CurranMoens:2002}
James~R. Curran and Marc Moens.
\newblock 2002.
\newblock Improvements in automatic thesaurus extraction.
\newblock In {\em Proceedings of the ACL-02 Workshop on Unsupervised Lexical
  Acquisition - Volume 9}, ULA '02, pages 59--66.

\bibitem[\protect\citename{Deerwester \bgroup et al.\egroup
  }1990]{Deerwesteretal:1990}
Scott~C. Deerwester, Susan~T. Dumais, Thomas~K. Landauer, George~W. Furnas, and
  Richard~A. Harshman.
\newblock 1990.
\newblock Indexing by latent semantic analysis.
\newblock {\em Journal of American Society for Information Science},
  41(6):391--407.

\bibitem[\protect\citename{{Delli Bovi} \bgroup et al.\egroup
  }2015]{dellibovi-espinosaanke-navigli:2015:EMNLP}
Claudio {Delli Bovi}, Luis {Espinosa Anke}, and Roberto Navigli.
\newblock 2015.
\newblock Knowledge base unification via sense embeddings and disambiguation.
\newblock In {\em Proceedings of the 2015 Conference on Empirical Methods in
  Natural Language Processing}, pages 726--736, Lisbon, Portugal, September.
  Association for Computational Linguistics.

\bibitem[\protect\citename{Erk}2007]{Erk:07}
Katrin Erk.
\newblock 2007.
\newblock A simple, similarity-based model for selectional preferences.
\newblock In {\em Proceedings of the 45th Annual Meeting of the Association for
  Computational Linguistics, { Prague, Czech Republic}}.

\bibitem[\protect\citename{Espinosa-Anke \bgroup et al.\egroup
  }2016]{espinosa2016extasem}
Luis Espinosa-Anke, Horacio Saggion, Francesco Ronzano, and Roberto Navigli.
\newblock 2016.
\newblock {ExTaSem! Extending, Taxonomizing and Semantifying Domain
  Terminologies}.
\newblock In {\em Proceedings of the 30th Conference on Artificial Intelligence
  (AAAI'16)}.

\bibitem[\protect\citename{Hassan and Mihalcea}2011]{HassanMihalcea:2011}
Samer Hassan and Rada Mihalcea.
\newblock 2011.
\newblock Semantic relatedness using salient semantic analysis.
\newblock In {\em Proceedings of AAAI}, pages 884,889.

\bibitem[\protect\citename{Huang \bgroup et al.\egroup }2012]{Huangetal:2012}
Eric~H. Huang, Richard Socher, Christopher~D. Manning, and Andrew~Y. Ng.
\newblock 2012.
\newblock Improving word representations via global context and multiple word
  prototypes.
\newblock In {\em Proceedings of ACL}, pages 873--882, Jeju Island, Korea.

\bibitem[\protect\citename{Iacobacci \bgroup et al.\egroup
  }2015]{iacobacci:2015}
Ignacio Iacobacci, Mohammad~Taher Pilehvar, and Roberto Navigli.
\newblock 2015.
\newblock {SensEmbed}: Learning sense embeddings for word and relational
  similarity.
\newblock In {\em Proceedings of ACL}, pages 95--105, Beijing, China.

\bibitem[\protect\citename{Jauhar \bgroup et al.\egroup
  }2015]{jauhar2015ontologically}
Sujay~Kumar Jauhar, Chris Dyer, and Eduard Hovy.
\newblock 2015.
\newblock Ontologically grounded multi-sense representation learning for
  semantic vector space models.
\newblock In {\em Proceedings of NAACL}.

\bibitem[\protect\citename{Johansson and Pi\~{n}a
  Nieto}2015]{johansson2015embedding}
Richard Johansson and Luis Pi\~{n}a Nieto.
\newblock 2015.
\newblock Embedding a semantic network in a word space.
\newblock In {\em Proceedings of NAACL}, pages 1428--1433.

\bibitem[\protect\citename{Jurgens and
  Pilehvar}2016]{jurgens-pilehvar:2016:SemEval}
David Jurgens and Mohammad~Taher Pilehvar.
\newblock 2016.
\newblock {SemEval-2016} task 14: Semantic taxonomy enrichment.
\newblock In {\em Proceedings of the 10th International Workshop on Semantic
  Evaluation}, pages 1092--1102, San Diego, California, June.

\bibitem[\protect\citename{Jurgens \bgroup et al.\egroup
  }2014]{jurgens2014semeval}
David Jurgens, Mohammad~Taher Pilehvar, and Roberto Navigli.
\newblock 2014.
\newblock Semeval-2014 task 3: Cross-level semantic similarity.
\newblock {\em SemEval 2014}, page~17.

\bibitem[\protect\citename{Laender \bgroup et al.\egroup
  }2002]{Laenderetal:2002}
Alberto H.~F. Laender, Berthier~A. Ribeiro-Neto, Altigran~S. da~Silva, and
  Juliana~S. Teixeira.
\newblock 2002.
\newblock A brief survey of web data extraction tools.
\newblock {\em SIGMOD Rec.}, 31(2):84--93.

\bibitem[\protect\citename{Li and Jurafsky}2015]{li-jurafsky:2015:EMNLP}
Jiwei Li and Dan Jurafsky.
\newblock 2015.
\newblock Do multi-sense embeddings improve natural language understanding?
\newblock In {\em Proceedings of the 2015 Conference on Empirical Methods in
  Natural Language Processing}, pages 1722--1732, Lisbon, Portugal, September.

\bibitem[\protect\citename{Liu \bgroup et al.\egroup }2015]{liu2015topical}
Yang Liu, Zhiyuan Liu, Tat-Seng Chua, and Maosong Sun.
\newblock 2015.
\newblock Topical word embeddings.
\newblock In {\em AAAI}, pages 2418--2424.

\bibitem[\protect\citename{Mikolov \bgroup et al.\egroup
  }2013a]{Mikolovetal:2013}
Tomas Mikolov, Kai Chen, Greg Corrado, and Jeffrey Dean.
\newblock 2013a.
\newblock Efficient estimation of word representations in vector space.
\newblock {\em CoRR}, abs/1301.3781.

\bibitem[\protect\citename{Mikolov \bgroup et al.\egroup
  }2013b]{mikolov2013distributed}
Tomas Mikolov, Ilya Sutskever, Kai Chen, Greg~S Corrado, and Jeff Dean.
\newblock 2013b.
\newblock Distributed representations of words and phrases and their
  compositionality.
\newblock In {\em Advances in neural information processing systems}, pages
  3111--3119.

\bibitem[\protect\citename{Miller \bgroup et al.\egroup }1990]{Milleretal:90}
George~A. Miller, R.T. Beckwith, Christiane~D. Fellbaum, D.~Gross, and
  K.~Miller.
\newblock 1990.
\newblock {W}ord{N}et: an online lexical database.
\newblock {\em International Journal of Lexicography}, 3(4):235--244.

\bibitem[\protect\citename{Moro \bgroup et al.\egroup }2014]{Moroetal:14tacl}
Andrea Moro, Alessandro Raganato, and Roberto Navigli.
\newblock 2014.
\newblock {Entity Linking meets Word Sense Disambiguation: a Unified Approach}.
\newblock {\em Transactions of the Association for Computational Linguistics
  (TACL)}, 2:231--244.

\bibitem[\protect\citename{Navigli and Ponzetto}2012]{NavigliPonzetto:12aij}
Roberto Navigli and Simone~Paolo Ponzetto.
\newblock 2012.
\newblock {B}abel{N}et: {T}he automatic construction, evaluation and
  application of a wide-coverage multilingual semantic network.
\newblock {\em Artificial Intelligence}, 193:217--250.

\bibitem[\protect\citename{Navigli}2009]{navigli:09}
Roberto Navigli.
\newblock 2009.
\newblock {W}ord {S}ense {D}isambiguation: {A} survey.
\newblock {\em ACM Computing Surveys}, 41(2):1--69.

\bibitem[\protect\citename{Neelakantan \bgroup et al.\egroup
  }2014]{Neelakantanetal:2014}
Arvind Neelakantan, Jeevan Shankar, Alexandre Passos, and Andrew McCallum.
\newblock 2014.
\newblock Efficient non-parametric estimation of multiple embeddings per word
  in vector space.
\newblock In {\em Proceedings of EMNLP}, pages 1059--1069, Doha, Qatar.

\bibitem[\protect\citename{Niemann and Gurevych}2011]{niemann2011people}
Elisabeth Niemann and Iryna Gurevych.
\newblock 2011.
\newblock {The people's web meets linguistic knowledge: automatic sense
  alignment of Wikipedia and Wordnet}.
\newblock In {\em Proceedings of the Ninth International Conference on
  Computational Semantics}, pages 205--214.

\bibitem[\protect\citename{Pantel and Lin}2002]{PantelLin:2002}
Patrick Pantel and Dekang Lin.
\newblock 2002.
\newblock Discovering word senses from text.
\newblock In {\em Proceedings of KDD}, pages 613--619.

\bibitem[\protect\citename{Pennacchiotti \bgroup et al.\egroup
  }2008]{Pennacchiottietal:2008}
Marco Pennacchiotti, Diego De~Cao, Roberto Basili, Danilo Croce, and Michael
  Roth.
\newblock 2008.
\newblock Automatic induction of {FrameNet} lexical units.
\newblock In {\em Proceedings of the Conference on Empirical Methods in Natural
  Language Processing}, EMNLP '08, pages 457--465.

\bibitem[\protect\citename{Pilehvar and Collier}2016]{PilehvarCollier:2016}
Mohammad~Taher Pilehvar and Nigel Collier.
\newblock 2016.
\newblock De-conflated semantic representations.
\newblock In {\em Proceedings of the Conference on Empirical Methods in Natural
  Language Processing}, Austin, TX, USA.

\bibitem[\protect\citename{Pilehvar and Navigli}2013]{PilehvarNavigli:2013}
Mohammad~Taher Pilehvar and Roberto Navigli.
\newblock 2013.
\newblock Paving the way to a large-scale pseudosense-annotated dataset.
\newblock In {\em Proceedings of NAACL-HLT}, Atlanta, USA.

\bibitem[\protect\citename{Pilehvar and Navigli}2014]{PilehvarNavigli:2014a}
Mohammad~Taher Pilehvar and Roberto Navigli.
\newblock 2014.
\newblock A robust approach to aligning heterogeneous lexical resources.
\newblock In {\em Proceedings of ACL}, pages 468--478.

\bibitem[\protect\citename{Pilehvar and Navigli}2015]{PilehvarNavigli:2015aij}
Mohammad~Taher Pilehvar and Roberto Navigli.
\newblock 2015.
\newblock From senses to texts: An all-in-one graph-based approach for
  measuring semantic similarity.
\newblock {\em Artificial Intelligence}, 228:95--128.

\bibitem[\protect\citename{Pilehvar \bgroup et al.\egroup
  }2013]{Pilehvaretal:2013}
Mohammad~Taher Pilehvar, David Jurgens, and Roberto Navigli.
\newblock 2013.
\newblock {A}lign, {D}isambiguate and {W}alk: a {U}nified {A}pproach for
  {M}easuring {S}emantic {S}imilarity.
\newblock In {\em Proceedings of the 51st Annual Meeting of the Association for
  Computational Linguistics}, pages 1341--1351, Sofia, Bulgaria.

\bibitem[\protect\citename{Radinsky \bgroup et al.\egroup }2011]{Radinsky:2011}
Kira Radinsky, Eugene Agichtein, Evgeniy Gabrilovich, and Shaul Markovitch.
\newblock 2011.
\newblock A word at a time: computing word relatedness using temporal semantic
  analysis.
\newblock In {\em Proceedings of WWW}, pages 337--346, Hyderabad, India.

\bibitem[\protect\citename{Reisinger and Mooney}2010]{ReisingerMooney:2010}
Joseph Reisinger and Raymond~J. Mooney.
\newblock 2010.
\newblock Multi-prototype vector-space models of word meaning.
\newblock In {\em Proceedings of ACL}, pages 109--117.

\bibitem[\protect\citename{Rothe and Sch\"utze}2015]{RotheSchutze:2015}
Sascha Rothe and Hinrich Sch\"utze.
\newblock 2015.
\newblock {AutoExtend: Extending Word Embeddings to Embeddings for Synsets and
  Lexemes}.
\newblock In {\em Proceedings of ACL}, pages 1793--1803, Beijing, China, July.
  Association for Computational Linguistics.

\bibitem[\protect\citename{Snow \bgroup et al.\egroup }2007]{Snowetal:2007}
Rion Snow, Sushant Prakash, Daniel Jurafsky, and Andrew~Y. Ng.
\newblock 2007.
\newblock Learning to merge word senses.
\newblock In {\em Proceedings of the 2007 Joint Conference on Empirical Methods
  in Natural Language Processing and Computational Natural Language Learning
  (EMNLP-CoNLL)}, pages 1005--1014, Prague, Czech Republic.

\bibitem[\protect\citename{{\v{S}}uster \bgroup et al.\egroup
  }2016]{vsuster2016bilingual}
Simon {\v{S}}uster, Ivan Titov, and Gertjan van Noord.
\newblock 2016.
\newblock Bilingual learning of multi-sense embeddings with discrete
  autoencoders.
\newblock In {\em Proceedings of NAACL-HLT}.

\bibitem[\protect\citename{Tian \bgroup et al.\egroup
  }2014]{tian2014probabilistic}
Fei Tian, Hanjun Dai, Jiang Bian, Bin Gao, Rui Zhang, Enhong Chen, and Tie-Yan
  Liu.
\newblock 2014.
\newblock A probabilistic model for learning multi-prototype word embeddings.
\newblock In {\em COLING}, pages 151--160.

\bibitem[\protect\citename{Turney and Pantel}2010]{TurneyPantel:10}
Peter~D. Turney and Patrick Pantel.
\newblock 2010.
\newblock From frequency to meaning: Vector space models of semantics.
\newblock {\em Journal of Artificial Intelligence Research}, 37:141--188.

\bibitem[\protect\citename{Turney \bgroup et al.\egroup }2003]{Turney:2003}
Peter~D. Turney, Michael~L. Littman, Jeffrey Bigham, and Victor Shnayder.
\newblock 2003.
\newblock Combining independent modules to solve multiple-choice synonym and
  analogy problems.
\newblock In {\em Proceedings of Recent Advances in Natural Language
  Processing}, pages 482--489, Borovets, Bulgaria.

\bibitem[\protect\citename{Vu and Parker}2016]{vu2016k}
Thuy Vu and D~Stott Parker.
\newblock 2016.
\newblock K-embeddings: Learning conceptual embeddings for words using context.
\newblock In {\em Proceedings of NAACL-HLT}, pages 1262--1267.

\bibitem[\protect\citename{Wu and Giles}2015]{wu2015sense}
Zhaohui Wu and C~Lee Giles.
\newblock 2015.
\newblock Sense-aaware semantic analysis: A multi-prototype word representation
  model using wikipedia.
\newblock In {\em AAAI}, pages 2188--2194. Citeseer.

\end{thebibliography}

\documentclass[11pt]{article}
\usepackage{acl2016}
\usepackage{times}
\usepackage{latexsym}
\usepackage{comment}
\usepackage{graphicx}
\usepackage{url}

\aclfinalcopy 

\newcommand{\GG}[1]{}

\title{Semantic Representations of Word Senses and Concepts}

\author{Jos\'e Camacho-Collados, Ignacio Iacobacci, Roberto Navigli \and Mohammad Taher Pilehvar$^\dagger$  \\
	    Department of Computer Science, Sapienza University of Rome\\
	      	 {\tt \{collados,iacobacci,navigli\}@di.uniroma1.it}\\
	$^\dagger$Language Technology Lab, DTAL, University of Cambridge\\
        {\tt     mp792@cam.ac.uk}}

\date{}

\begin{document}

\maketitle

\section{Introduction}

 Representing the semantics of linguistic items in a machine-interpretable form has been a major goal of Natural Language Processing since its earliest days. Among the range of different linguistic items, words have attracted the most research attention. However, word representations have an important limitation: they conflate different meanings of a word into a single vector. Representations of word senses have the potential to overcome this inherent limitation. Indeed, the representation of individual word senses and concepts has recently gained in popularity with several experimental results showing that a considerable performance improvement can be achieved across different NLP applications upon moving from word level to the deeper sense and concept levels. Another interesting point regarding the representation of concepts and word senses is that these models can be seamlessly applied to other linguistic items, such as words, phrases and sentences.

This tutorial\footnote{Slides available at \url{http://goo.gl/az7tBD}} will first provide a brief overview of the recent literature concerning word representation (both count and neural network based). It will then describe the advantages of moving from the word level to the deeper level of word senses and concepts, providing an extensive review of state-of-the-art systems. Approaches covered will not only include those which draw upon knowledge resources such as WordNet, Wikipedia, BabelNet or FreeBase as reference, but also the so-called multi-prototype approaches which learn sense distinctions by using different clustering techniques. Our tutorial will discuss the advantages and potential limitations of all approaches, showing their most successful applications to date. We will conclude by presenting current open problems and lines of future work.

\section{Outline}

\subsection{Semantic Representation: Foundations}

This session provides the necessary background for semantic representation. We will briefly cover the traditional vector space model \cite{TurneyPantel:10} followed by the more recent approaches based on neural networks \cite{Mikolovetal:2013}. We then provide reasons for the need to produce semantic representations for the deeper word sense level, focusing on the main limitation of the word-based approaches which is their inherent ambiguity. Finally, we show how sense-based representations are bound to overcome these limitations, hence providing improvements across several tasks.

\subsection{Knowledge-based sense representations}
    
We start this session by briefly introducing some of the most popular lexical knowledge resources that have been used by different sense representation techniques. We put emphasis on WordNet \cite{Milleretal:90}, the de facto standard sense inventory in the community, and Wikipedia, the largest collaboratively-constructed resource of the type, both of which have been extensively used by many researchers in the area. We discuss the advantages each of these resources provides and show how they are usually viewed as semantic networks and exploited for representation purposes.

Then, we provide a deep review of different techniques that learn representations for individual concepts in a target sense inventory. We cover all the existing approaches that model concepts in WordNet \cite{PilehvarNavigli:2015aij}, articles in Wikipedia \cite{HassanMihalcea:2011}, or concepts in larger sense inventories such as BabelNet \cite{iacobacci:2015,camachocollados:2016aij} or FreeBase  \cite{bordes2011learning,bordes2013translating}. We will also cover some approaches that make use of additional external corpora (or word representations learned on the basis of statistical clues) besides the target knowledge resource \cite{chenunified:2014,chen2015improving,johansson2015embedding,jauhar2015ontologically,RotheSchutze:2015,PilehvarCollier:2016}. We discuss the advantages of these knowledge-based representations and focus on how neural network-based learning has played a role in this area in the past few years.

\subsection{Unsupervised sense representations}

In this session we cover the so-called multi-prototype techniques that learn multiple representations per word, each corresponding to a specific meaning of the word. We will illustrate how these approaches leverage clustering algorithms for dividing the contexts of a word into multiple contexts for its different meanings \cite{ReisingerMooney:2010,Huangetal:2012,Neelakantanetal:2014,tian2014probabilistic,wu2015sense,li-jurafsky:2015:EMNLP,liu2015topical,vu2016k,vsuster2016bilingual}.

\subsection{Advantages and limitations}

This session reviews some of the advantages and limitations of the knowledge-based and unsupervised techniques, describing the applications for which they are suitable and mentioning some issues such as the knowledge acquisition bottleneck.

\subsection{Applications}

This session focuses on different applications of sense representations. We briefly mention some of the main applications and tasks to which sense representations can be applied. Sense representations may be used in virtually every task in which word representations have been traditionally applied. Examples of such tasks include automatic thesaurus generation \cite{Crouch:1988,CurranMoens:2002}, information extraction \cite{Laenderetal:2002}, semantic role labelling \cite{Erk:07,Pennacchiottietal:2008}, and word similarity \cite{Deerwesteretal:1990,Turney:2003,Radinsky:2011,mikolov2013distributed} and clustering \cite{PantelLin:2002}. We will provide comparisons between word and sense representations performance, discussing the advantages and limitations of each approach. Moreover, we will show how sense representations can also be applied to a wide variety of additional tasks such as entity linking and word sense disambiguation \cite{navigli:09,chenunified:2014,camachocolladosetal:2015b,RotheSchutze:2015,camachocollados:lrec2016}, sense clustering \cite{Snowetal:2007,camachocolladosetal:2015}, alignment of lexical resources \cite{niemann2011people,PilehvarNavigli:2014a}, taxonomy learning \cite{espinosa2016extasem}, knowledge-base completion \cite{bordes2013translating}, information extraction \cite{dellibovi-espinosaanke-navigli:2015:EMNLP},  or sense-based semantic similarity \cite{BudanitskyHirst:06,Pilehvaretal:2013,iacobacci:2015}, to name a few.

\subsection{Open problems and future work}

This last session provides a summary of possible directions of future work on semantic sense representation. We discuss various problems associated with the current representation approaches and propose lines of research in order to effectively apply sense representations in natural language understanding tasks.

\section{Instructors}

\paragraph{Jos\'e Camacho Collados} is a Google Doctoral Fellow and PhD student at the Sapienza University of Rome\footnote{\url{http://wwwusers.di.uniroma1.it/~collados/}}, working under the supervision of Prof. Roberto Navigli. His research focuses on Natural Language Processing and on the area of lexical semantics in particular. He has developed \textsc{Nasari}\footnote{\url{http://lcl.uniroma1.it/nasari/}} \cite{camachocollados:2016aij}, a novel semantic vector representation for concepts based on Wikipedia that features a high coverage of named entities and has been successfully used on different NLP tasks. Jos\'e has also worked on the development of evaluation benchmarks for word and sense representations \cite{camacho2015framework,camachocolladosrepeval:2016} and is the co-organizer of a SemEval 2017 task on multilingual and cross-lingual semantic similarity. His background education includes an Erasmus Mundus Master in Natural Language Processing and Human Language Technology and a 5-year BSc degree in Mathematics.

\paragraph{Ignacio Iacobacci} is a PhD student at the Sapienza University of Rome\footnote{\url{https://iiacobac.wordpress.com/}}, working under the supervision of Prof. Roberto Navigli. His research interests lie in the fields of Machine Learning, Natural Language Processing, Neural Networks. He is currently working on Word Sense Disambiguation and Distributional Semantics. Ignacio presented SensEmbed\footnote{\url{http://lcl.uniroma1.it/sensembed/}} at ACL 2015 \cite{iacobacci:2015}, a novel approach for word and relational similarity built from exploiting semantic knowledge for modeling arbitrary word senses in a large sense inventory. His background includes a MSc. in Computer Science and 8 years as a developer including 4 years as a Machine Learning - NLP specialist.

\paragraph{Roberto Navigli} is an Associate Professor in the Department of Computer Science at La Sapienza University of Rome and a member of the Linguistic Computing Laboratory\footnote{\url{http://wwwusers.di.uniroma1.it/~navigli/}}. His research interests lie in the field of Natural Language Processing, including: Word Sense Disambiguation and Induction, Ontology Learning, Knowledge Representation and Acquisition, and multilinguality. In 2007 he received a Ph.D. in Computer Science from La Sapienza and he was awarded the Marco Cadoli 2007 AI*IA national prize for the Best Ph.D. Thesis in Artificial Intelligence. In 2013 he received the Marco Somalvico AI*IA prize, awarded every two years to the best young Italian researcher in Artificial Intelligence. He is the creator and founder of BabelNet\footnote{\url{http://www.babelnet.org}} \cite{NavigliPonzetto:12aij}, both a multilingual encyclopedic dictionary and a semantic network, and its related project Babelfy\footnote{\url{http://www.babelfy.org}} \cite{Moroetal:14tacl}, a state-of-the-art multilingual disambiguation and entity linking system. He is also the Principal Investigator of MultiJEDI\footnote{\url{http://multijedi.org/}}, a 1.3M euro 5-year Starting Grant funded by the European Research Council and the responsible person of the Sapienza unit in LIDER, an EU project on content analytics and language technologies. Moreover, he is the Co-PI of "Language Understanding cum Knowledge Yield" (LUcKY), a Google Focused Research Award on Natural Language Understanding.

\paragraph{Mohammad Taher Pilehvar} is a Research Associate in the Language Technology Lab of the University of Cambridge\footnote{\url{http://www.pilevar.com/taher/}} where he is currently working on NLP in the biomedical domain. Taher completed his PhD in 2015 under the supervision of Prof. Roberto Navigli. Taher's research lies in lexical semantics, mainly focusing on semantic representation, semantic similarity, and Word Sense Disambiguation. He has co-organized three SemEval tasks \cite{jurgens2014semeval,jurgens-pilehvar:2016:SemEval} and has authored multiple conference and journal papers on semantic representation and similarity in top tier venues. He is the first author of a paper on semantic similarity that was nominated for the best paper award at ACL 2013 \cite{PilehvarNavigli:2013}.

\section*{Acknowledgments}
The authors gratefully acknowledge the support of the ERC Starting
  Grant MultiJEDI No.\ 259234.

\bibliography{acl2016}
\bibliographystyle{acl2016}

\end{document}